OpenStudy (anonymous):

rationalize the denominator and simplify. (3+√3)/(3-√3)

OpenStudy (ja1):

I don't get this math >_<

OpenStudy (preetha):

Music, I don't know how to do this. But I pinged @hero and he is very smart.

hero (hero):

Multiply the numerator and denominator by (3+√3), then simplify. When it says to 'rationalize the denominator', it means to multiply top and bottom by its conjugate. The conjugate of (3-√3) is (3+√3). Therefore: \[\frac{3 + \sqrt{3}}{3 - \sqrt{3}} \times \frac{3 + \sqrt{3}}{3 + \sqrt{3}}\] @music101, can you simplify that?

OpenStudy (anonymous):

yeah that's the part I know how to do, Hero

OpenStudy (anonymous):

so....1?

OpenStudy (anonymous):

that's the answer

OpenStudy (anonymous):

thank u hero!

OpenStudy (ja1):

hero is the best and welcome to OpenStudy @music101

OpenStudy (anonymous):

thanks ☺

OpenStudy (anonymous):

the problem is.... 1 is not an option!!

hero (hero):

It's not an option because that's not quite what the correct simplification. Do you have a way of showing what you did?

OpenStudy (anonymous):

ugh! dang it! I wish I did. everything's on paper though

hero (hero):

I'm interested in knowing how you got 1. My result isn't close to that.

hero (hero):

Can you try using the draw button to write it out?

OpenStudy (anonymous):

yes. hold on let me double-check something

OpenStudy (anonymous):

wait... I was looking at the wrong problem :p

OpenStudy (anonymous):

my final answer is 2+√3???

hero (hero):

That's correct.

OpenStudy (anonymous):

ok thanks ☺

hero (hero):

Great recovery

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

the next problem says to solve for x: √(4x-2) -1=2

OpenStudy (anonymous):

I got x= 11/4

OpenStudy (anonymous):

that's correct rite?

hero (hero):

Correct again

OpenStudy (preetha):

Thanks Hero!

OpenStudy (anonymous):

i'm a little stuck on this problem it's pretty easy but Idk where i'm wrong can u tell me if my answer is rite?

hero (hero):

Which one?

OpenStudy (anonymous):

(√(4-3x)/(√3x)=2

OpenStudy (anonymous):

I got x= 4/15

OpenStudy (anonymous):

& what abt this? x+2=√(4-x)

hero (hero):

\[\frac{\sqrt{4 - 3x}}{\sqrt{3x}} = 2\]\[\sqrt\frac{{4 - 3x}}{{3x}} = 2\]\[\frac{{4 - 3x}}{{3x}} = 4\]\[4 - 3x = 4(3x)\]\[4 - 3x = 12x\]\[4 = 15x\]\[\frac{15}{4} = x\]

hero (hero):

I think you just simplified backwards a bit.

OpenStudy (anonymous):

ohh :( but I submitted it as my final answer

OpenStudy (anonymous):

ok well.....

hero (hero):

Whoops, you had it right all along.

hero (hero):

4/15 = x is correct.

OpenStudy (anonymous):

oh... ok then

hero (hero):

\[(x + 2) = \sqrt{4 - x}\] \[(x + 2)^2 = 4 - x\] Expand \((x + 2)^2\)

OpenStudy (anonymous):

x+2=√(4-x) solve for x?

OpenStudy (anonymous):

how do u know which x to solve for?

OpenStudy (anonymous):

so.. x^2+4x+4 rite?

hero (hero):

Okay good so you have \[x^2 + 4x + 4 = 4 - x\] Anytime you see that, it means to move everything to one side and create a quadratic equation.

hero (hero):

@music101, did you go offline?

OpenStudy (anonymous):

sorry i'm back